# reference.bib

### v2

@@ 2.2 Solving the forward problem
In order to solve (2.6) we must be able to solve (2.1) in an efficient and stable way. In the past, many methods have been developed: Eulerian methods [filbet2003comparison], Lagrangian methods [Verboncoeur2005] and, semi-Lagrangian methods [cheng1976integration]. In this work we will use the latter mainly because they are fully explicit and unconditionally stable (they do not suffer from a CFL condition).
In a nutshell, semi-Lagrangian methods trace the characteristics exactly and perform an interpolation since the translated solution may not coincide with the grid. The literature proposes different ways to perform this interpolation, with spline based [cheng1976integration, sonnendrucker1999semi, filbet2003comparison] and Fourier-based methods [Klimas1994, Klimas2018] being the most famous. More recently, the interpolation has been done following a discontinuous Galerkin approach [Crouseilles2011, qiu2011positivity, Rossmanith2011, einkemmer2019performance, Einkemmer2024]. In our case, we deploy a simple linear interpolation.
@@
$\left\{\begin{array}[]{rcl}E_{\mathsf{f}^{\ast}}&=&\partial_{x}V_{\mathsf{f}^{\ast}}\\ -\partial_{xx}V_{\mathsf{f}^{\ast}}&=&\rho_{\text{p}}\end{array}\right.\,.$ (2.8)
@@
$\left\{\begin{array}[]{rcl}\hat{E}_{\mathsf{f}^{\ast}}&=&2\pi i\xi\hat{V}_{\mathsf{f}^{\ast}}\\ 4\pi^{2}\xi^{2}\hat{V}_{\mathsf{f}^{\ast}}&=&\hat{\rho}_{\text{p}}\end{array}\right.\,.$
Since we can compute $\rho_{\text{p}}$ by knowing $f_{\text{eq}}$ and, $\mathsf{f}^{\ast}$ from the previous step, we have that
$\hat{E}_{\mathsf{f}^{\ast}}(t,\xi)=-\frac{1}{2\pi i\xi}\hat{\rho}_{\text{p}}(t,\xi)\,.$
@@
We summarize the semi-Lagrangian method described above in Algorithm $1$https://github.com/maguerrap/Vlasov-Poisson.
@@
$\{x_{i},v_{j}\}_{i,\,j=0}^{M-1}$, $f_{\text{iv}}$, $\Delta t$, $N$, $f_{\text{eq}}$, $L_{x}$, $L_{v}$, $H$.
2:$\bm{f}_{n}:=\{\mathsf{f}(t^{n},x_{i},v_{j})\}_{i,\,j=0}^{M-1}$ for $n=0,...,N$.
3:for $n=0,1,...,N$ do
4:     Compute $\mathsf{f}^{\ast}(x_{i},v_{j})=\mathsf{f}(t^{n},x_{i}-v_{j}\Delta t/2,v_{j})$, for $i,j=0,1,...,M-1$.
@@
6:     Compute $\mathsf{f}^{\ast\ast}(x_{i},v_{j}+(E_{\mathsf{f}^{\ast}}+H)\Delta t)$, for $i,j=0,1,...,M-1$.
7:     Compute $\mathsf{f}(t^{n+1},x_{i},v_{j})=\mathsf{f}^{\ast\ast}(x_{i}-v_{j}\Delta t/2,v_{j})$, for $i,j=0,1,...,M-1$.
@@ 2.3.1 Two Stream example
As a first example, we consider the Two Stream equilibrium. This distribution arises when two streams (or beams) of charged particles moving relative to each other are interpenetrating, causing instability. Mathematically, the equilibrium distribution is given by
@@
$f_{\text{iv}}(x,v)=\left(1+\varepsilon\cos(k_{0}x)\right)f_{\text{eq}}(v)\,,$ (2.10)
with $\varepsilon=0.001$.
@@
Figure 1: Simulation of (2.1) with $H\equiv 0$ for the Two Stream equilibrium. From left to right we have $f_{\text{eq}}$, $f(T=30,x,v)$, $E_{f}(t,x)$, $\mathcal{E}_{f}(t)$.
@@ 2.3.2 Bump-on-Tail example
The Bump-on-Tail equilibrium is another steady state that is heavily investigated. This distribution appears when a small group of fast-moving electrons (a “bump”) is superimposed on the tail of the background electron velocity distribution and this can cause instability. This can happen due to particle injection or wave-particle interactions. The equilibrium is defined as
@@
$f_{\text{iv}}(x,v)=f_{\text{eq}}(v)+\frac{\sqrt{2}\,\varepsilon}{10\sqrt{\pi}}\exp\left(-2(v-\bar{v}_{2})^{2}\right)\cos(0.1x)\,,$ (2.12)
with $\varepsilon=0.001$, the electric energy will exponential grow in time. In Figure 2 we simulate (2.1) using Algorithm 1 without an external electric field (i.e., $H\equiv 0$). The solution demonstrates strong electric energy growth at about $T=40$ when we terminate the simulation. The simulation is conducted over a $L_{x}\times L_{v}=20\pi\times 9.0$ box.
Figure 2: Simulation of (2.1) with $H\equiv 0$ for the Bump-on-Tail equilibrium. From left to right we have $f_{\text{eq}}$, $f(T=40,x,v)$, $E_{f}(t,x)$, $\mathcal{E}_{f}(t)$.
@@ 3 Dispersion relation and linear stability analysis
Examining the dispersion relation is a classical tool to investigate a dynamical system’s stability features. Mathematically, this amounts to conducting linearization of the system around a desired equilibrium state, and examining the PDE spectrum of the resulting linear PDE operator. If the PDE operator has positive spectrum, the system is expected to demonstrate exponential growth. However, the extra control term $H$ can shift the PDE spectrum and the desired control is to find the $H$ so that the positive spectrum are eliminated.
This strategy was proposed and studied in depth in [Einkemmer2025], and for the completeness of the current paper, we briefly summarize the results. These analytical findings will shed lights to our simulation.
We first linearize the system around the desired equilibrium state, writing $f=f_{\text{eq}}+f_{\text{p}}$, then omitting the higher order terms in the expansion, the leading order solution to $f_{\text{p}}$ is:
$&\partial_{t}f_{\text{p}}(t,x,v)+v\partial_{x}f_{\text{p}}(t,x,v)-\left[E(t,x)+H(t,x)\right]\partial_{v}f_{\text{eq}}(v)=0,\\ &\partial_{xx}V(t,x)=-\rho_{\text{p}}(t,x),\\ &E(t,x)=\partial_{x}V(t,x)\,.\end{aligned}\right.$ (3.1)
We define three crucial transforms:
@@
We can deduce a very important identity:
@@
According to Mellin inversion Laplace transform formula, $\rho_{\text{p}}$ will have exponential growth if $L[\hat{\rho}_{\text{p}}]$ has singularity on $\mathcal{R}(s)>0$, the right half of the complex plane for $s$. This would be the case if $01+L[\hat{U}(\cdot,k)](s)=$ 0 for some $(k,s)$ while the whole right hand side does not vanish at the some point. Consequently, to eliminate the exponential growth, one should also impose a root for the right hand side when the factor $1+L[\hat{U}(\cdot,k)](s)$ touches zero. In light of this discussion, in [Einkemmer2025], the authors propose to set the control $H$ according to:
$-L[\hat{S}(\cdot,k)](s)+ikL[\hat{H}(\cdot,k)](s)=h\left(1+L[\hat{U}(\cdot,k)](s)\right),$ (3.6)
@@
In particular,
  • By setting $h(\cdot)=0$, the control $H$ becomes
    $ik\hat{H}(t,k)=\hat{S}(t,k)\,,$
    and the solution is a free-streaming:
    $\rho_{\text{p}}(t,x)=S(t,x)=\int f_{\text{p}}(t,x-vt,v)\,\mathrm{d}v\,.$
  • By setting $h(\cdot)=\alpha\times\cdot$, we have:
    $ik\hat{H}(t,k)=\hat{S}(t,k)+\alpha\left(\delta(t)+\hat{U}(t,k)\right),$ (3.8)
    and the density solution is:
    ${\rho}_{\text{p}}(t,x)={S}(t,x)+\alpha{U}(t,x)\,.$ (3.9)
    where $U$ is the inverse Fourier transform of $\hat{U}$.
It is evident that the control $H$ designed above has both time and space dependence. However, the $H$ we are seeking for only has spatial dependence, rendering these analysis results not immediately useful. However, the study does provide a nice range to make a guess for optimization. In particular, in both designs for $H$ above, roughly $ik\hat{H}\approx\hat{S}$. So whenever $\hat{S}(k,\cdot)\equiv 0$ for a fixed wavenumber $k$, meaning that the initial perturbation lacks the mode corresponding to this frequency, it is therefore appropriate to set $\hat{H}(k)=0$.
Taking the two-stream example shown in Section 2.3.1,
$f_{\text{p}}(t=0)=\epsilon\cos(k_{0}x)f_{\text{eq}}\quad\Rightarrow\quad\rho_{\text{p}}(t=0)=\epsilon\cos(k_{0}x)\,,$
where $f_{\text{eq}}$ is defined in (2.9). So a good guess for $H$ should simply be:
$H(x)=\mathcal{O}({\epsilon\left(\sin(k_{0}x)+\cos(k_{0}x)\right)})\,.$
The parameters still need to be fine tuned. The analysis above, however, pins an important frequency for the control ($k_{0}$), and the important scaling ($\epsilon$).
@@ 4 Landscape analysis of the objective function
We look into landscape analysis for the objective function in this section. This is to examine the dependence of the three objective functions (defined in ($L^{2}$)-(KL)-(EE)) on parameters of the control (defined in (2.5)). Landscape analysis requires a sweep of PDE solves over all parameter choices in a pre-defined domain, and thus is computationally demanding. We are only presenting results for the following two choices of parameter settings:

### v3

@@ 2.2 Solving the forward problem
In order to solve (2.6) we must be able to solve (2.1) in an efficient and stable way. In the past, many methods have been developed: Eulerian methods [21], Lagrangian methods [51] and, semi-Lagrangian methods [14, 45]. A general summary can be found in [22]. In this work we will use the latter mainly because they are fully explicit and unconditionally stable (they do not suffer from a CFL condition [5, 11]).
In a nutshell, semi-Lagrangian methods trace the characteristics exactly and perform an interpolation since the translated solution may not coincide with the grid. The literature proposes different ways to perform this interpolation, with spline based [14, 48, 21] and Fourier-based methods [32, 33] being the most famous. More recently, the interpolation has been done following a discontinuous Galerkin approach [15, 44, 45, 17, 19]. In our case, we deploy a simple linear interpolation.
@@
$\left\{\begin{array}[]{rcl}E_{\mathsf{f}^{\ast}}&=&\partial_{x}V_{\mathsf{f}^{\ast}}\\ \partial_{xx}V_{\mathsf{f}^{\ast}}&=&1-\rho_{\mathsf{f}^{\ast}}\end{array}\right.\,$. (2.8)
@@
$\left\{\begin{array}[]{rcl}\hat{E}_{\mathsf{f}^{\ast}}&=&2\pi i\xi\hat{V}_{\mathsf{f}^{\ast}}\\ 4\pi^{2}\xi^{2}\hat{V}_{\mathsf{f}^{\ast}}&=&\hat{\rho}_{\mathsf{f}^{\ast}}\,,\quad(\forall\xi\neq 0)\end{array}\right.\,$.
Since we can compute $\hat{\rho}_{\mathsf{f}^{\ast}}$ by knowing $\mathsf{f}^{\ast}$ from the previous step, we have that
$\hat{E}_{\mathsf{f}^{\ast}}(t,\xi)=-\frac{1}{2\pi i\xi}\hat{\rho}_{\mathsf{f}^{\ast}}(t,\xi)\,.$
@@
We summarize the semi-Lagrangian method described above in Algorithm 1222https://github.com/maguerrap/Vlasov-Poisson.
@@
[…]
3:for n=0,1,,Nn=0,1,...,N do
4:  Compute $\mathsf{f}^{\ast}(x_{i},v_{j})=\mathsf{f}(t^{n},x_{i}-v_{j}\Delta t/2,v_{j})$, for i,j=0,1,,M1i,j=0,1,...,M-1.
@@
6:  Compute $\mathsf{f}^{\ast\ast}(x_{i},v_{j}+(E_{\mathsf{f}^{\ast}}+H)\Delta t)$, for i,j=0,1,,M1i,j=0,1,...,M-1.
7:  Compute $\mathsf{f}(t^{n+1},x_{i},v_{j})=\mathsf{f}^{\ast\ast}(x_{i}-v_{j}\Delta t/2,v_{j})$, for i,j=0,1,,M1i,j=0,1,...,M-1.
@@ 2.3.1 Two Stream example
As a first example, we consider the Two Stream equilibrium. This distribution arises when two streams (or beams) of charged particles moving relative to each other are interpenetrating, causing instability [3]. Mathematically, the equilibrium distribution is given by
@@
$f_{\text{iv}}(x,v)=\left(1+\varepsilon\cos(k_{0}x)\right)f_{\text{eq}}(v)\doteq f_{\text{eq}}(v)+f_{\text{p}}(t=0,x,v)\,$, (2.10)
with the initial perturbation $f_{\text{p}}(t=0,x,v)$ being of order $\varepsilon=0.001$.
@@
Figure 1: Simulation of (2.1) with $H\equiv 0$ for the Two Stream equilibrium. From left to right we have $f_{\text{eq}}(v)$, $f(T=30,x,v)$, $E_{f}(t,x)$ and $\mathcal{E}_{f}(t)$.
@@ 2.3.2 Bump-on-Tail example
The Bump-on-Tail equilibrium is another steady state that is heavily investigated. This distribution appears when a small group of fast-moving electrons (a “bump”) is superimposed on the tail of the background electron velocity distribution and this can cause instability [42]. This can happen due to particle injection or wave-particle interactions. The equilibrium is defined as
@@
$f_{\text{iv}}(x,v)=f_{\text{eq}}(v)+\underbrace{\frac{\sqrt{2}\,\varepsilon}{10\sqrt{\pi}}\exp\left(-2(v-\bar{v}_{2})^{2}\right)\cos(0.1x)}_{\doteq f_{\text{p}}(t=0,x,v)}\,$, (2.12)
with the initial perturbation $f_{\text{p}}(t=0,x,v)$ being of order $\varepsilon=0.001$, the electric energy will exponential grow in time. In Figure 2 we simulate (2.1) using Algorithm 1 without an external electric field (i.e., $H\equiv 0$). The solution demonstrates strong electric energy growth at about $T=40$ when we terminate the simulation. The simulation is conducted over a $L_{x}\times L_{v}=20\pi\times 9.0$ box.
Figure 2: Simulation of (2.1) with $H\equiv 0$ for the Bump-on-Tail equilibrium. From left to right we have $f_{\text{eq}}(v)$, $f(T=40,x,v)$, $E_{f}(t,x)$ and $\mathcal{E}_{f}(t)$.
@@ 3 Dispersion relation and linear stability analysis
Examining the dispersion relation is a classical approach to investigating the stability properties of dynamical systems [39]. Mathematically, this involves linearizing the system around a desired equilibrium state and analyzing the spectrum of the resulting linear PDE operator. If the operator has spectrum with positive real parts, the system exhibits exponential divergence from equilibrium. To counteract this, we introduce a control term $H$ aimed at modifying the spectrum—specifically, to eliminate its unstable (positive) components and thereby keep the system close to the desired state.
This strategy was proposed and studied in depth in [18], where the dispersion relation was derived both with and without an external field, and a general procedure was outlined to design a field capable of suppressing growth. However, the method targets complete suppression by eliminating all unstable roots, which is a stringent requirement. Achieving this demands highly flexible external fields—indeed, the designs in [18] result in fields that oscillate significantly in time. While theoretically sound, such time-dependent fields are physically impractical and incompatible with our assumption (2.5).
In contrast, we seek to impose a far less flexible external field: one that contains only finitely many Fourier modes and is time-independent. This constraint necessitates a relaxation of the original goal. Rather than eliminating all unstable roots, we shift our focus to suppressing the modes that grow the fastest. Our approach begins by identifying the specific modes that would exhibit the strongest exponential growth in the absence of control, then targeting their associated roots for elimination. This analysis builds on our previous derivation of the dispersion relation, now adapted to account for time-independent control.
Ultimately, the system remains nonlinear, so previously unexcited modes may be triggered over time. As a result, the present strategy is effective only in the short term. For long-term suppression, a return to a full PDE-constrained optimization framework is necessary. Nonetheless, our analysis offers a good initial guess for such optimization-based methods.
In the following section, we begin by reviewing the dispersion relation derived in [18], as presented in subsection 3.1. We then extend the analysis to the case of a time-independent external field, and propose our strategy in subsection 3.2. It provides a good initial guess for further optimization.
3.1 Dispersion relation and control strategy when $H=H(t,x)$
We firstly briefly review the dispersion relation for the Vlasov–Poisson system subject to an external field $H(t,x)$ that varies in both time and space. The original derivation is complete in [18].
To do so, we first linearize the system around the desired equilibrium state, writing $f=f_{\text{eq}}+f_{\text{p}}$, then omitting the higher order terms in the expansion, the leading order solution to $f_{\text{p}}$ is:
$\left\{\begin{aligned} &\partial_{t}f_{\text{p}}(t,x,v)+v\partial_{x}f_{\text{p}}(t,x,v)-\left[E_{f_{\text{p}}}(t,x)+H(t,x)\right]\partial_{v}f_{\text{eq}}(v)=0\,,\\ &E_{f_{\text{p}}}(t,x)=\partial_{x}V_{f_{\text{p}}}(t,x)\,,\\ &\partial_{xx}V_{f_{\text{p}}}(t,x)=-\rho_{\text{p}}(t,x)=-\int f_{\text{p}}(t,x,v)\,\mathrm{d}v\,.\end{aligned}\right$. (3.1)
Three crucial transforms will be heavily leaned on:
@@ 3.1 Dispersion relation and control strategy when $H=H(t,x)$
Repeatedly deploying these transforms, one can deduce an important identity:
@@
According to Mellin inversion Laplace transform formula, $\rho_{\text{p}}$ will have exponential growth if $L[\hat{\rho}_{\text{p}}]$ has a singularity on $\Re(s)>0$, while the source term on the right hand side of (3.3) does not vanish at the same roots. As a consequence, to eliminate the exponential growth, one should also force the right hand side to be zero at the roots of $1+L[\hat{U}(\cdot,k)](s)$. In light of this discussion, in [18], the authors propose to set the control $H$ according to:
$L[\hat{S}(\cdot,k)](s)+ikL[\hat{H}(\cdot,k)](s)=h\left(1+L[\hat{U}(\cdot,k)](s)\right),$ (3.6)
@@
This design effectively eliminates the positive roots, and thus the PDE’s instability is suppressed. Further analytical derivations and comprehensive numerical validations are provided in [18].
3.2 Dispersion relation and control strategy when $H=H(x)$
It is evident that the control $H$ designed above has both time and space dependence. However, the $H$ we are seeking for only has spatial dependence, rendering these analysis results not immediately useful. In this section, we modify the analysis above and re-derive the dispersion relation with $H$ assumed to be time-independent.
Similar to the derivation above, we apply the Fourier–Laplace transform and find:
$\bigl{(}L[\hat{\rho}_{\mathrm{p}}(\cdot,k)](s)-L[\hat{S}(\cdot,k)](s)\bigr{)}\,\bigl{(}1+L[\hat{U}(\cdot,k)](s)\bigr{)}=L[\hat{U}(\cdot,k)](s)\,\bigl{(}-L[\hat{S}(\cdot,k)](s)+\tfrac{ik\,\hat{H}(k)}{s}\bigr{)}.$ (3.8)
In the derivation, we used $L[1](s)=\frac{1}{s}$ and the fact that $H$ does not have time dependence.
Following the strategy as in (3.6), we should set
[…]
for all complex frequencies $s$. The equation is pointwise in $(k,s)$ while the function to be sought for only varies in $k$, rendering no solution that satisfies the entire equation. However, as discussed, in this regime when we have a far less flexible external field, we ought to relax our original goal. In this particular setting, we are to find the fastest growing mode and aim at suppressing that mode only.
Mathematically, to find the fastest growing mode, we set
$s_{0}(k)=\text{argmax}\{\Re(s)>0:\;1+L\bigl{[}\hat{U}(\cdot,k)\bigr{]}(s)=0\}$ (3.9)
for every fixed $k$. This is to identify the Laplace index who are roots for the operator and has the largest real component. We then require $\hat{H}$ only to eliminate that specific component: $L\bigl{[}\hat{S}(\cdot,k)\bigr{]}(s_{0}(k))\;+\;\frac{ik\,\hat{H}(k)}{s_{0}}=0$. This yields
$\hat{H}(k)\;=\;\frac{s_{0}(k)\,L\bigl{[}\hat{S}(\cdot,k)\bigr{]}(s_{0}(k))}{ik}\,.$ (3.10)
We summarize our strategy as follows. For a given equilibrium distribution $U$ and the initial perturbation $S$, we:
  1. 1.
    compute (3.9) for $s_{0}(k)$ for all $k$ seen in the initial perturbation;
  2. 2.
    compute the spectral amplitude $\hat{H}(k)$ using (3.10).
Since $\hat{S}$ is determined, according to (3.5), by the initial perturbation. By definition (2.10) and (2.12), it is of order $\mathcal{O}(\varepsilon)$, hence the guessed control $H$ should also be of $\mathcal{O}(\varepsilon)$.
Example 3.1.
We apply this strategy on Two Stream instability and Bump-on-Tail instability.
  • Two Stream instability: The initial perturbation (2.10) only sees $k=\pm 1$, so we only compute the associated $s_{0}(\pm 1)$. As shown in Figure 3 (left panel) that plots $|1+L(\hat{U}(\cdot,1))|(s)$, we find that the numerical root that has the biggest real component is
    $s_{0}(\pm 1)=0.236\,.$
    Substituting it into (3.10), we have:
    $\hat{H}(1)=-3.15768\times 10^{-9}+0.000643706\,i,\quad\hat{H}(-1)=-3.15768\times 10^{-9}-0.000643706\,i,$
    and thus
    $H(x)=-0.00128741\,\sin(k_{0}x)-6.31537\times 10^{-9}\,\cos(k_{0}x)$
  • Bump-on-tail instability: The initial perturbation (2.12) only sees $k=\pm 1$, so we only compute the associated $s_{0}(\pm 1)$. As shown in Figure 3 (right panel) that plots $|1+L(\hat{U}(\cdot,1))|(s)$, we find that the numerical root that has the biggest real component is
    $s_{0}(\pm 1)=0.230\mp 0.324\,i\,.$
    Substituting it into (3.10), we have:
    $\hat{H}(1)=7.05072\times 10^{-4}+1.60134\times 10^{-4}\,i,\quad\hat{H}(-1)=7.05072\times 10^{-4}-1.60134\times 10^{-4}\,i.$
    and thus
    $H(x)=-0.000320268\,\sin(k_{0}x)+0.00141014\,\cos(k_{0}x)\,.$
Figure 3: Norm of $\|1+\mathcal{L}[U](\cdot,1)(s)\|$: the Two Stream (left) and the Bump-on-Tail (right) examples. In Two Stream, the minimum occurs at $s_{0}=0.236+0i$ with $\|1+\mathcal{L}[U](\cdot,1)(s_{0})\|=9.564044\times 10^{-4}$. In Bump-on-Tail, the minimum occurs at $s_{0}=0.230-0.324i$ with $\|1+\mathcal{L}[U](\cdot,1)(s_{0})\|=3.756278\times 10^{-3}$.
These calculations yield a strong initial guess for the external field $H$. When applied to equation (2.1), we observe significant suppression effects, as illustrated in Figure 4. However, it is important to note that the electric energy eventually exhibits exponential growth over long times. This reflects the onset of nonlinear dynamics, which cannot be fully addressed by our initial design. As such, complete suppression requires the use of a PDE-constrained optimization framework. The field $H$ constructed here serves as an effective initial guess for such optimization, which we will discuss in detail in Section 5.
Figure 4: Simulation of (2.1) for the Two Stream equilibrium up to $T=30$ (top) and Bump-on-Tail equilibrium up to $T=40$ (bottom). In each row, from left to right: $f_{\text{eq}}(v)$, $f(T,x,v)$ (for $H\equiv 0$), $f(T,x,v)$ (for $H$ obtained from Example 3.1) and $\mathcal{E}_{f}(t)$ (for $H\equiv 0$ and $H$ obtained from Example 3.1).
@@ 4 Landscape analysis of the objective function
In this section, we examine the landscape of the objective function. This is to evaluate the dependence of the four objective functions (defined in (KL), (EE), (KLT) and (EET)) on parameters of the control (defined in (2.5)). Landscape analysis requires a sweep of PDE solves over all parameter choices in a pre-defined domain, and thus is computationally demanding.
In Section 4.1 and Section 4.2 respectively, we provide the landscape for choices of two different parameter settings for the two different canonical examples (Sections 2.3.1 and 2.3.2) respectively.
4.1 Two Stream example
We report results for the Two Stream instability problem in this section. Here we have the following two choices of parameter settings: